OpenStudy (anonymous):

which class has 28% boys

OpenStudy (anonymous):

I need more information.

OpenStudy (anonymous):

answer choices are: A. Class A: 3 out of 20 B. Class B: 28 out of 40 C. Class C: 1 out of 28 D. Class D: 14 out of 50

OpenStudy (anonymous):

How do you determine the percentage of something?

OpenStudy (anonymous):

i dont know that is why i need help

OpenStudy (anonymous):

The percentage of something is another way of representing a fraction. The percent sign "%" means per 100. So if you have 25%, that means that is 25/100. 77% 77/100 and so on. The formula usually given to find percentage is: \[\frac{x}{n} \times 100\] Where n is the total number in a group and x is the subset you are trying to find the percentage of. Example: I take a quiz. There are 36 possible points on the quiz. I get 27 points. The percentage on my quiz is calculated like this: \[\frac{27}{36} \times 100 = 75\%\]